OpenStudy (phantomcrow):

How would you find the area bounded between y^2 -x = 2 and x -y = 0

OpenStudy (phantomcrow):

I don't know what should be my upper bound...

OpenStudy (agent0smith):

First thing you need is a diagram. The upper and lower bounds will come from the intersections of the curves.

OpenStudy (phantomcrow):

Would it then be possible to find where those curves intersect and use those are your bounds?

OpenStudy (phantomcrow):

Based of, say, a system instead of graphing?

OpenStudy (phantomcrow):

Nevermind, I got it.

OpenStudy (agent0smith):

Any time you are asked to find area or volume in a calculus question, you should use a diagram.

OpenStudy (sshayer):

\[y^2=x+2\] It is a parabola with vertex at (-2,0) x-y=0 is a straight line. x=y \[x^2-x-2=0\] \[x^2-2x+x-2=0\] \[x(x-2)+1(x-2)=0\] \[(x+1)(x-2)=0,\] x=-1,2 as x=y point of intersections are (-1,-1) and (2,2)

OpenStudy (sshayer):